OpenStudy (anonymous):

One yard company charges $10x + $50 a month to mow and edge a lawn. A second company charges $8x + $60 a month to mow and edge a lawn. Each lawn company also offers additional services for an extra fee. Let x represent the number of additional services. Mr. Mead is comparing the two companies. How many additional services are needed for the two companies to be the same price?

OpenStudy (anonymous):

set the two equations equal to each other and solve for x 10x +50 = 8x +60 -50 -50 ----------------- 10x = 8x + 10 10x = 8x + 10 -8x -8x --------------- 2x = 10 2x = 10 divide by 2 on both sides x=5 five additional services are needed for the two companies prices to be equal